DGS Maintainance Day Confirmation

History

 No problem

 No problem

 Pico Ping:
  ETMY (rebooted on 2022/1/6)
  PR2_IM: NG => reboot.

 No problem

 Pico Ping:
  SR2_IM: NG => reboot.

 stepper:
  ETMX_F1_1: Abnormality disappeared as of last week (Nov.25 10:32:50 JST).
 Pico Ping:
  SR2_IM: NG => reboot.

 Pico Ping:
  SR3_IM, ITMX: NG => reboot

 No problem

 No problem

 Pico Ping
  SR3(BF), BS(IM):NG => reboot

 Pico Ping
  SR3(BF), BS(IM):NG => reboot

 No problem

 No problem

 Stepper:
  BS_F0Y LimitSwitch: NG
 Pico Ping:
  ETMY: NG (Network Switch Down?)
  SR2_IM, SR2_STM: NG 
 Temperature:
  ETMY: NG (Network Switch Down?)

 Pico Ping:
  SR3_IM: NG
  ETMX: NG
  ETMY: NG
  SR2_STM: NG

 Pico Ping:
  BS_BF:NG
  SR2_STM:NG

 Pico Ping:
  ETMX: NG
  SR2(IM): NG

KAGRA/Subgroups/VIS/OpsManual/DGSMaintainance (last edited 2022-01-21 15:36:52 by satoru.ikeda)